# Laminator: Verifiable ML Property Cards using Hardware-assisted Attestations

### v1

@@ V-A The Prover
We implement Laminator’s prover with the aid of the Gramine framework [37], which allows the execution of unmodified applications inside an SGX enclave. This allows Laminator to produce attestations of properties computed using off-the-shelf libraries such as PyTorch [51].
We implement the attestations from Section IV. The different inputs and outputs from different attestation enclaves are shown in Figure 3. The dataset enclave takes $\mathcal{D}_{tr}$ as an input and generates the distribution along with DistAtt. The training enclave takes the $\mathcal{D}_{tr}$, $\mathcal{M}_{ar}$, and $\mathcal{T}$ as input, to train a model inside the enclave, and output the trained model $\mathcal{M}$ and PoT. Given $\mathcal{M}$ and $\mathcal{D}_{te}$, the metric enclave can generate different metrics for accuracy, fairness, and robustness, along with their corresponding attestations as an output. These metrics are used to evaluate $\mathcal{M}$ before deployment. Finally, once $\mathcal{M}$ is deployed, it can be run inside an inference enclave such that for each input, we get the corresponding $\mathcal{M}$’s output along with IOAtt.
@@
SGX EnclavesTraining Dataset($\mathcal{D}^{prov}_{tr}$)Training and ModelConfiguration ($\mathcal{T}$+$\mathcal{M}_{ar}$)Dataset Enclave….Training Enclave….DistributionDistAttModelPoTInputInference Enclave….Metric Enclave….Test Dataset($\mathcal{D}_{te}$)MetricAccAttFairAttRobAttOutputIOAtt
@@
Figure 3: Inputs and outputs for different attestation enclaves. Laminator includes four different enclaves which are indicated in orange. Attestations are in ellipse, outputs are in blue, and inputs to enclaves are white. We indicate measurer in green.
@@
The attestations are generated from a custom ML library built atop PyTorch. A hash of the Python code responsible for these attestations, along with PyTorch and all other files used by the enclave, is embedded into the enclave using the Gramine framework, which provides a library OS that allows the integrity of the Python code to be verified when read from disk, incorporating it into the MRENCLAVE value included in SGX attestations.
@@
Laminator includes a wrapper for Python’s open library call, which is used to read input files into memory, hashing them at the same time, allowing PyTorch to randomly access measured files without risk of time-of-check-time-of-use (TOCTOU) vulnerabilities. We implement an additional mechanism to check the resulting hash against a manifest, allowing for a single short measurement of large directories of input files without the need to keep them all in memory.
@@
After the computation is complete, a hash of the output is incorporated into the user data field of an SGX report, which is then used to obtain a DCAP quote [3] that can be verified by remote parties. This output takes the form of a JSON string representing a fragment of the model card metadata [^3] https://github.com/huggingface/hub-docs/blob/main/modelcard.md, where a model is named with the hash of its file.
V-B The Verifier
@@
  • A set of serialized attestations in JSON format, each containing an SGX quote—representing the attestation signature from Section IV—and base64-encoded output.
@@
$\mathcal{V}$ validates each quote, validating it against the output of the respective attestation, then using the enclave measurement from the quote to look up the respective certification, which contains a JSON structure that serves as a template for the payload, and which is checked according the following rules:
  • null in the certification matches anything in the payload
  • Dictionaries in the certification match dictionaries (or arrays of dictionaries) in the payload whose keys match those of the certification, and whose values match the values of the certification
  • Strings, booleans, and integers in the certification match identical values (or arrays of identical values) in the payload
  • Other values are not allowed in the certification
@@
For each model referenced by an attestation, $\mathcal{V}$ produces a model card, formed by taking the union of the various model card metadata fragments to produce a single JSON dictionary that is reformatted into the YAML contents of the model card model-index field, which contains experimental data in a structured form.
V-C Authorization
@@
Enclaves must only generate attestations of properties that they are not certified for and $\mathcal{V}$ should not trust an unknown enclave or trust an attestation from an enclave which it is not certified to make. To check for this, an endorser can provide an external certificate for an enclave, in addition to those for datasets or models, to indicate which properties each enclave is allowed to generate. These external certificates can be included in the assertion bundle and validated during the verification, though in our proof-of-concept they are hardcoded into the verifier.
VI Evaluation
We present the experimental setup which includes datasets, model architecture, training configuration, and metrics (Section VI-A). We then present a performance evaluation of different attestations (Section VI-B).
VI-A Experimental Setup
@@
We now present different datasets, model architecture and training configuration for each of the dataset, and metrics for evaluation.
Datasets. We consider two datasets, CIFAR and CENSUS, for evaluating different ML property attestations which will help design verifiable ML property cards. Both the datasets are standard benchmark datasets used in most of the ML literature. We describe these datasets below.
CIFAR consists of 60000 images of 32 $\times$ 32 colour images belonging to one of ten classes of different objects: airplane, bird, cat, deer, automobile, dog, frog, horse, ship, and truck. The dataset contains 6000 images per class. We use 50000 training images and 10000 test images.
CENSUS consists of 48842 data records with 42 attributes about individuals from the 1994 US Census data obtained from UCI Machine Learning dataset repository. It includes sensitive attributes such as sex and race of the participant. Other attributes include marital status, education, occupation, job hours per week among others. The classification task is to estimate whether the individual’s annual income is at least 50,000 USD. We used 24,000 training data records and 24,000 testing data records. This dataset has two sensitive attributes, race and sex, which we use for attesting distributional properties and fairness.
Model Architecture. We now describe the model architecture used for both the datasets. For CIFAR, we use a convolutional neural network (CNN) model with three convolution layers of dimensions: [32, 64, 128]. Each being a filter with padding of 1 and kernel size of 3x3. After each convolution layer, there is a maxpool layer of size of 2x2 and a ReLU activation function and a batch normalization layer. The final layer has 512 nodes with a dropout layer with 0.25 probability of dropout. For CENSUS, we use a multi-layered perceptron (MLP) with three layers with each hidden layer of dimensions: [32, 64, 32]. Further, we use Tanh activation function.
Training Configuration. We train the model on CIFAR with 52 epochs with a learning rate of 1e-3, using Adam optimizer, and a batch size of 64 achieving accuracy of 81.31%. For CENSUS, we train the model with 10 epochs, learning rate of 1e-3, Adam optimizer, and batch size of 256 achieving accuracy 84.43%. We chose the number of epochs based on when the accuracy converges.
Metrics. We evaluate Laminator, breaking down the computation into four phases:
  • Startup which includes time spent starting the Python runtime, and in the case of Laminator, initializing the Gramine-based SGX enclave. This is a one-time-per-enclave-instance cost that is amortizable over multiple attestation requests.
  • Data pre-processing which includes time spent reading the data from disk, as well as performing any measurements or preprocessing.
  • Computation which includes the time spent on the inference, training, or property measurement operation.
  • Attestation which includes the time spent performing the attestation.
VI-B Performance Evaluation
We present the performance numbers for different attestations across the three ML property cards.
Attestations for Datasheets. We measure the performance of two attestations: distribution attestation, and proof of training.
@@ VI-B Performance Evaluation
Distribution Attestation. We evaluate distributional properties only for CENSUS dataset since it has sensitive attributes, race and sex. We generate attestations indicating i) the distribution of a sensitive attribute $z$ and ii) the distribution of $z$ given the classification label $y$ (denoted $z|y$ in Table II) We present the performance results for distributional properties of both race and sex attributes in Table II.
TABLE II: Performance of distributional property attestation.
Operation CENSUS-Race CENSUS-Sex
$z$ $z|y$ $z$ $z|y$
Startup 80 $\pm$ 9 s 80 $\pm$ 7 s 93 $\pm$ 11 s 75 $\pm$ 0.44 s
Preprocessing 330 $\pm$ 23 ms 320 $\pm$ 3.9 ms 370 $\pm$ 4.2 ms 320 $\pm$ 1.7 ms
Computation 14 $\pm$ 0.5 ms 27 $\pm$ 0.16 ms 15 $\pm$ 0.11 ms 27 $\pm$ 0.16 ms
Attestation 35 $\pm$ 71 ms 5.9 $\pm$ 0.15 ms 6.5 $\pm$ 0.55 ms 5.8 $\pm$ 0.079 ms
We find that the time for computation and attestation are low indicating good efficiency regardless of whether the sensitive attribute is race or sex. This is observed as the lack of significant difference in the time required to compute $z$ and $z|y$ of CENSUS-Race and CENSUS-Sex. Further, as expected, the computation time under the column for $z$ is less than $z|y$, due to the additional conditioning on the true classification labels $y$.
Attestations for Model Cards. We measure the performance of two attestations for verifiable model cards: proof of training, and accuracy attestations. We leave the evaluation of robustness and fairness attestations as they are similar to accuracy attestation but using different metrics.
Proof of Training. We train models for CIFAR and CENSUS inside Intel SGX and measure the time taken. We present the time averaged over ten runs indicated in milliseconds in Table III.
TABLE III: Performance of proof of training.
Operation CIFAR CENSUS
Startup 81 $\pm$ 0.82 s 71 $\pm$ 0.2 s
Preprocessing 4400 $\pm$ 24 ms 1300 $\pm$ 7 ms
Computation 3060 $\pm$ 120 s 36 $\pm$ 3.9 s
Attestation 5.4 $\pm$ 0.15 ms 5.8 $\pm$ 0.13 ms
We show that training an entire model for CIFAR takes $\sim$51 minutes while CENSUS takes $\sim$6 minutes. This is significantly more efficient than proof of training using ZKPs which is limited to simple models and datasets [17].
Accuracy Attestation. We measure the performance for both CIFAR and CENSUS datasets with 10,000 and 24,000 data records respectively. We present the results in Table IV.
TABLE IV: Performance of accuracy attestation.
Operation CIFAR (ms) CENSUS (ms)
Startup 36 $\pm$ 0.17 s 71 $\pm$ 0.17 s
Preprocessing 290 $\pm$ 2 ms 1300 $\pm$ 8 ms
Computation 3500 $\pm$ 100 ms 120 $\pm$ 3 ms
Attestation 5.7 $\pm$ 0.2 ms 5.7 $\pm$ 0.02 ms
The results show that the computation time for CIFAR is higher than CENSUS which is expected given the size of each CIFAR image is 32$\times$32$\times$3 compared to tabular data record with 42 attributes. Despite this, it takes only 3.5s to compute test accuracy on CIFAR while 0.1s for CENSUS.
@@
Input-Model-Output Attestation. We evaluate the performance of input-model-output attestation by averaging 10 runs of 100 inferences. We present these results in Table V. For 100 inferences, the computation only takes 71ms for CIFAR10 compared to 21ms for CENSUS.
TABLE V: Performance of input-model-output attestation.
Operation CIFAR CENSUS
Startup 32 $\pm$ 0.19 s 71 $\pm$ 0.19 s
Preprocessing 0.50 $\pm$ 1.7 ms 6.3 $\pm$ 0.10 ms
Computation 71 $\pm$ 41 ms 21 $\pm$ 0.40 ms
Attestation 6.6 $\pm$ 1.1 ms 5.9 $\pm$ 0.20 ms
Summary. These findings highlight the efficiency of Laminator, with minimal overhead incurred during attestation generation. Notably, this initial startup cost is followed by the ability for limitless verification by any number of parties, affirming the scalability of our approach. Furthermore, TEEs exhibit versatility in training various models, exemplified by their successful implementation with complex architectures like AlexNet. In contrast, zero-knowledge proofs, constrained by model-dependent proofs, are restricted to simpler models.
VI-C Revisiting Requirements
We now indicate how the Laminator satisfies different requirements from Section III-B.
Efficiency. Running models inside SGX to generate attestations incurs a low overhead as, instead of developing custom ML pipeline, we can use high performance libraries like Pytorch along with the Intel’s AMX extensions. As shown above, $\mathcal{P}$ has some initial overhead but once started, multiple attestations can be generated with minimal overhead.
Once the attestation is made, the cost of verification is low, involving only certificate checking, and does not depend on the size of the model. Hence, Laminator satisfies R1.
Scalability. For each prover, attestation is done once and then can be verified independently by multiple verifiers. This process is non-interactive, and so the complexity of generating and verifying the quote does not depend on either the size of the model or the number of verifiers. Finally, the assertion bundle can combine many different attestations from different enclaves, which are then converted into a standard format; this means that Laminator can scale to large numbers of independent provers. All these design choices result in Laminator satisfying R2.
Versatility. As Laminator’s properties are measured for attestation purely in software, while hardware being used only to isolate this software, it is easy to extend Laminator to new properties. A developer can write a new Python script measuring the desired property. The enclave is built and its measurement certified, e.g., by a trusted endorser, or by hardcoding it into the verifier by its developer, enabling $\mathcal{V}$ to ensure that the new enclave can be trusted to attest its claimed assertions. This is aided by the attestation and certification formats, which can be directly mapped to YAML-formatted property card metadata, meaning that $\mathcal{V}$ need not be modified to introduce new metrics or ML property cards. Hence, Laminator satisfies R3.
VII Related Work

### v3

@@ 5.1. The Prover
We implement Laminator’s prover using Gramine (Kuvaiskii et al., 2022), which allows unmodified applications to run inside an SGX enclave. This allows Laminator to generate attestations for properties computed with off-the-shelf libraries like PyTorch (Paszke et al., 2019).
We implement the attestations from Section 4. The different inputs and outputs from different attestation enclaves are shown in Figure 3:
  1. (1)
    the dataset enclave takes $\mathcal{D}_{tr}$ as an input and generates the distribution along with Algorithm 2.
  2. (2)
    the training enclave takes $\mathcal{D}_{tr}$, $\mathcal{M}_{ar}$, and $\mathcal{T}$ as input, to train a model inside the enclave, and output the trained model $\mathcal{M}$ along with PoT (Algorithm 3).
  3. (3)
    the metric enclave takes $\mathcal{M}$ and $\mathcal{D}_{te}$ as input, generating metrics for accuracy, fairness, and robustness, along with their attestations as an output (Algorithm 4, Algorithm 5, and Algorithm 6 respectively).
  4. (4)
    the inference enclave runs $\mathcal{M}$ after deployment, producing the model’s output and corresponding IOAtt for each input (Algorithm 7).
@@
Trusted CertifierSEV-SNP Virtual MachineSGX EnclaveTraining Dataset($\mathcal{D}^{prov}_{tr}$)Training and ModelConfiguration ($\mathcal{T}$+$\mathcal{M}_{ar}$)Dataset Enclave….Training Enclave….DistributionDistAttModelPoTInputInference Enclave….Metric Enclave….Test Dataset($\mathcal{D}_{te}$)MetricAccAttFairAttRobAttOutputIOAtt
@@
Figure 3. Inputs and outputs for different attestation enclaves. Laminator includes four different enclaves which are indicated in orange. Attestations are in ellipse, outputs are in blue, and inputs to enclaves are white. We indicate measurer in green.
@@
The attestations are generated from a custom ML library built atop PyTorch. A hash of the Python code responsible for these attestations, along with PyTorch and all other files used by the enclave, is embedded into the enclave using Gramine, which provides a library OS that allows the integrity of the Python code to be verified when read from disk, incorporating it into the MRENCLAVE value included in SGX attestations.
@@
Laminator includes a wrapper for Python’s open library call, which is used to read input files into memory, hashing them at the same time, allowing PyTorch to randomly access measured files without risk of time-of-check-time-of-use (TOCTOU) vulnerabilities. We implement an additional mechanism to check the resulting hash against a manifest, allowing for a single short measurement to represent a large directory of input files without the need to keep them all in memory.
@@
After the computation is complete, a hash of the output is incorporated into the user data field of an SGX report, which is then used to obtain a DCAP quote (Intel, 2019) that can be verified by remote parties. This output takes the form of a JSON string representing a fragment of the model card metadata [^4] https://github.com/huggingface/hub-docs/blob/main/modelcard.md, where a model is named with the hash of its file.
5.2. The Verifier
@@
  • A set of serialized attestations in JSON format, each containing an SGX quote—representing the attestation signature from Section 4—and base64-encoded output.
@@
$\mathcal{V}$ checks each quote by validating it against the output of the corresponding attestation. $\mathcal{V}$ then uses the enclave measurement from the quote to look up the associated certification, which contains a JSON structure that acts as a template for the payload. This template is then validated according to the following rules:
  • null in the certification matches anything in the payload,
  • The dictionaries in the certification correspond to dictionaries (or arrays of dictionaries) in the payload. The keys in the payload must match those in the certification, and the values in the payload must match the values in the certification,
  • Strings, booleans, and integers in the certification match identical values (or arrays of identical values) in the payload,
  • Other values are not allowed in the certification.
@@
For each model referenced by an attestation, $\mathcal{V}$ produces a model card, formed by taking the union of the various model card metadata fragments to produce a single JSON dictionary that is reformatted into the YAML contents of the model card model-index field, which contains experimental data in a structured form. Also, Laminator lets $\mathcal{V}$ combine TEE attestations with other guarantees (e.g., external certificates in Section 4.5). Assurances from ZKP-like mechanisms can be integrated into Laminator similarly.
5.3. Authorization
@@
Enclaves should only generate attestations for certified properties, and the verifier should not trust attestations from unknown enclaves or those not certified to make a specific assertion. To check for this, an endorser can provide an external certificate for an enclave, in addition to those for datasets or models, to indicate which properties each enclave is allowed to generate. These external certificates can be included in the assertion bundle and validated during the verification, though in our proof-of-concept they are hardcoded into the verifier.
6. Evaluation
We now evaluate Laminator and our proof-of-concept against the requirements identified in Section 3.
6.1. R1: Efficiency
We assess efficiency by measuring the overhead from measurement and attestation in a confidential ML workload, exploring how it varies with model complexity and dataset size. We aim to validate the following conjectures: the overhead of Laminator is
  1. (1)
    determined by a roughly constant attestation overhead, and a measurement overhead dominated by the size of its inputs and outputs. We indicate the inputs and outputs for measurements across the four attestation types below:
    • Distribution attestation: Training dataset as input, and distribution value as output.
    • Proof of training: Training dataset as input, and trained model as output.
    • Accuracy attestation: Test dataset and trained model as input, and accuracy as output.
    • Input-model-output attestation: Trained model and data record for inference as input, and predictions as output.
  2. (2)
    small in comparison with the computational task being attested for all attestations except for IO attestation which we discuss separately in Section 6.1.2 (Input-Model-Output Attestation).
6.1.1. Experimental Setup
@@
We now present hardware configuration, datasets, model architecture and training configurations, and metrics for evaluation. Reported values are averages of five runs.
Hardware Configuration. We conduct an evaluation on a machine with the following components:
  • CPU: Intel Xeon Silver 4514Y.
  • RAM: 512GB, 8x 64GB (Micron Technology MTC40F2046S1RC48BA1 MHCC), 4800 MHz.
  • Storage: Micron 7450 MTFDKBG1T9TFR, 1920GB
  • Mainboard: Supermicro X13SEDW-F
Datasets. We consider three datasets of varying size drawn from several domains: a tabular dataset CENSUS (Kohavi, 1996), an image dataset UTKFACE (Zhang et al., 2017), and a textual dataset IMDB (Maas et al., 2011). These datasets are standard benchmark used in ML literature.
CENSUS contains 48,842 records with 42 attributes from the 1994 US Census, including sensitive attributes like sex and race. The task is to predict whether an individual’s annual income exceeds 50,000 USD. A 50-50 train-test split yields 24,000 records each for training and testing. This translates to 2MB disk size for train and test dataset.
UTKFACE is a large-scale face dataset with 48x48 RGB images for classifying individuals as young (¡30 years). It includes 11,852 training and 11,853 testing samples, with sensitive attributes of race and sex. This translates to 100MB disk size for train and test dataset. IMDB is a text dataset containing movie reviews from the IMDB website for binary classification of positive/negative sentiments. We use 37,500 training and 12,500 test data records. This translates to 49.5MB and 16.5MB disk size for train and test dataset respectively.
Model Architecture. We train two models on each dataset: a smaller model (tagged with the suffix “-S”), and a larger model (tagged with the suffix “-S”). We summarize the models and the number of parameters in Table 2.
Table 2. Summary of models across different datasets.
Model Description #Parameters Model Size (MB)
CENSUS-S MLP: [128] 12,290 0.05
CENSUS-L MLP: [128, 256, 512, 256] 308,482 1.2
UTKFACE-S VGG11 9,227,010 36.95
UTKFACE-L VGG16 14,724,162 58.96
IMDB-S LSTM: [64, 256, 256] 920,385 3.69
IMDB-L LSTM: [64, 256, 256, 256, 256] 1,973,057 7.60
CENSUS-S and CENSUS-L are multi-layer perceptron (MLP) models with hidden layers of dimensions [128] and [128, 256, 512, 256] respectively with $\tanh$ activation functions. CENSUS-S has $\sim$12K model parameters (0.05 MB) while CENSUS-L has $\sim$308K parameters (1.2 MB). We train these models for 5 epochs which gives an accuracy of 78.33% (CENSUS-S) and 81.09% (CENSUS-L).
UTKFACE-S and UTKFACE-L are VGG11 and VGG13 models (Simonyan, 2014) respectively. VGG11 with hidden layer dimensions: [64, ‘M’, 128, ‘M’, 256, 256, ‘M’, 512, 512, ‘M’, 512, 512, ‘M’] and VGG16 includes [64, 64, ‘M’, 128, 128, ‘M’, 256, 256, 256, ‘M’, 512, 512, 512, ‘M’, 512, 512, 512, ‘M’], where ‘M’ indicates maxpool. UTKFACE-S has $\sim$9.22 million parameters (36.95MB) and UTKFACE-L with $\sim$14.72 million parameters (58.96MB). We train the models for 10 epochs to get 80.28% (UTKFACE-S) and 80.81% (UTKFACE-L) accuracy.
IMDB-S and IMDB-L are LSTM models with hidden layer dimensions of [64, 256, 256] and [64, 256, 256, 256, 256] respectively. IMDB-S has $\sim$920K parameters (3.69MB) and IMDB-L has 1.97 million parameters (7.60MB). We train these models for five epochs to get an accuracy of 85.60% (IMDB-S) and 86.18% (IMDB-L).
Metrics. We split the tasks into the following phases of computation, with Laminator’s overhead marked in bold:
  • Input loading is the time spent reading the operation’s input data.
  • Input measurement is the time spent hashing the input data to produce an input measurement.
  • Data pre-processing is the time spent on transforming the inputs (e.g., feature selection, data augmentation).
  • Computation is the time spent on the inference, training, or property measurement.
  • Output measurement is the time spent hashing the output data to produce an output measurement.
  • Output storage is the time spent formatting and writing the output data.
  • Attestation is the time spent producing the attestation of the measurements.
Laminator’s overhead is therefore the total of input measurement, output measurement, and attestation. We use the total time spent on the remaining phases as the baseline, which includes input loading, data pre-processing, computation, and output storage, against which we compare Laminator’s overhead.
6.1.2. Results
We present the overheads of attesting four types of ML properties for datasheets, model cards, and inference cards.
Attestations for Datasheets. We measure the execution time of distributional property attestation for verifiable datasheets.
@@ 6.1.2. Results
Distribution Attestation. Table 3 summarizes the evaluation of distributional properties for CENSUS and UTKFACE, which include race and sex as sensitive attributes. We do not evaluate on the IMDB dataset since its text data does not contain any sensitive attributes. We generate attestations for two types of distributional properties: i) the distribution of a sensitive attribute $z$ and ii) the distribution of $z$ given the classification label $y$ (denoted $z|y$ in Table 3).
Table 3. Performance of distributional property attestation. Operations in gray form the baseline (orange), while those in blue form the Laminator’s overhead (orange). We omit ”Output Size” as the output is the distribution value, which is a single value of negligible size.
Dataset CENSUS UTKFACE
Preprocessing 9.90 s 54.22 s
Input Load 28.89 ms 1.78 s
Computation (Race: z) 73.29 ms 66.28 ms
Computation (Sex: z) 4.38 ms 3.78 ms
Computation (Race: z—y) 26.35 ms 23.49 ms
Computation (Sex: z—y) 10.43 ms 8.74 ms
Output Storage 1.05 ms 1.11 ms
Baseline 10.01 s 54.33 s
Input Meas. 5.90 ms 1.10 s
Input Size 2 MB 100 MB
Output Meas. 0.72 ms 0.87 ms
Attestation 29.85 ms 31.23 ms
Overhead 36.46 ms 1.14 s
Overhead (%) 0.36 2.05
Attestation time remains nearly constant across datasets. As expected, the input measurement time is higher for the larger datasets (UTKFACE). The total overhead incurred by Laminator varies between 0.36% and 2.05%. The size of the output is negligible so there will be some minimal output measurement time which is incurred (0.72 ms for CENSUS and 0.87 ms for UTKFACE).
Attestations for Model Cards. We report the overhead for attesting two types of ML card properties: proof of training and accuracy.
Proof of Training. We train both small and large models on CENSUS, UTKFACE, and IMDB (Table 4). The highest recorded overhead was 0.32% for the smallest model and dataset (CENSUS), while for larger datasets (UTKFACE and IMDB), Laminator ’s overhead is negligible, at 0.01% or lower. The attestation time does not increase, varying between 15ms and 31ms across each run, while the input measurement and output measurement times increase with the size of the input and output, as conjectured.
Table 4. Performance of proof of training. Operations in gray form the baseline (orange), while those in blue form the Laminator’s overhead (orange).
Dataset CENSUS UTKFACE IMDB
Model variant -S -L -S -L -S -L
Preprocessing 9.95 s 9.71 s 54.62 s 54.42 s 70.31 s 70.35 s
Input Load 26.49 ms 28.05 ms 1.80 s 1.83 s 0.33 s 0.35 s
Computation 973.24 ms 0.03 min 310.11 min 643.13 min 1803.16 min 3563.11 min
Output Storage 0.74 ms 0.74 ms 1.81 ms 1.84 ms 0.97 ms 1.09 ms
Baseline 10.95 s 11.53 s 311.05 min 644.06 min 1804.34 min 3564.29 min
Input Meas. 5.86 ms 5.88 ms 1.13 s 1.14 s 96.40 ms 97.35 ms
Input Size (MB) 2.00 2.00 100.00 100.00 49.50 49.50
Output Meas. (ms) 0.23 1.12 24.77 226.79 2.76 5.47
Output Size (MB) 0.05 1.20 36.95 58.96 3.69 7.60
Attestation (ms) 29.16 16.91 25.82 29.78 26.66 19.47
Overhead 35.25 ms 23.91 ms 1.18 s 1.39 s 125.82 ms 122.30 ms
Overhead (%) 0.32 0.21 0.01 0.01 0.00 0.00
Accuracy Attestation. We measure the accuracy attestation performance for both small and large model variants across all datasets (Table 5). The overhead of Laminator is less than 0.35%. The input measurement overhead increases with input size, while the attestation time varies between approximately between 22ms and 27ms, consistent with the other attestations. The size of the output is negligible so there will be some minimal output measurement time which is incurred, where the overhead for -S and -L models are close to each other. We note that there is a higher attestation time for -L models compared to -S models for CENSUS and UTKFACE. Since attestation is independent of the model size, we attribute this to some randomness during attestation.
Table 5. Performance of accuracy attestation. Operations in gray form the baseline (orange), while those in blue form the Laminator’s overhead (orange). We omit ”Output Size” as the output is the accuracy, which is a single value of negligible size.
Dataset CENSUS UTKFACE IMDB
Model variant -S -L -S -L -S -L
Preprocessing 9.46 s 10.14 s 54.74 s 55.15 s 70.17 s 70.37 s
Input Load 28.15 ms 34.34 ms 1.96 s 2.08 s 0.36 s 0.37 s
Computation 228.62 ms 306.03 ms 10.45 min 21.38 min 53.78 min 107.46 min
Output Storage 0.95 ms 0.90 ms 1.60 ms 1.51 ms 1.70 ms 1.74 ms
Baseline 9.72 s 10.48 s 11.40 min 22.33 min 54.95 min 108.64 min
Input Meas. 7.03 ms 8.68 ms 1.15 s 1.24 s 0.10 s 0.11 s
Input Size (MB) 2.05 3.20 136.95 158.96 20.19 24.10
Output Meas. 0.01 ms 0.02 ms 0.40 ms 0.32 ms 0.44 ms 0.33 ms
Attestation 26.92 ms 24.15 ms 25.93 ms 22.19 ms 22.61 ms 27.20 ms
Overhead 33.97 ms 32.85 ms 1.18 s 1.27 s 0.13 s 0.14 s
Overhead (%) 0.35 0.31 0.17 0.19 0.00 0.00
@@
Input-Model-Output Attestation. We measure the execution time of input-model-output attestation for a single inference averaged over 100 runs (Table 6). Since the size of the output is negligible the time for output measurement time is very small. Unlike the types of operations we have considered above, the baseline cost for a single inference is also very small. Consequently, it is no surprise that although Laminator’s computation time overhead for the attestation itself remains small (as in previous cases), the proportional overhead is relatively high, (between 39% and 3955%). The high overhead stems from the attestation signature being much more costly than inference computation, as shown in Table 6.
Table 6. Performance of input-model-output attestation. Operations in gray form the baseline (orange), while those in blue form the Laminator’s overhead (orange). We omit ”Output Size” as the output is the prediction, which is a single value of negligible size. We report the overhead computed using attestation as “w/ att” and that with signing using keypair as “w/ sgn”.
Dataset CENSUS UTKFACE IMDB
Model variant -S -L -S -L -S -L
Input Load 0.25 ms 6.13 ms 221.98 ms 336.42 ms 22.62 ms 41.47 ms
Computation 0.15 ms 0.40 ms 5.79 ms 14.05 ms 29.96 ms 57.97 ms
Output Storage 0.35 ms 1.33 ms 1.62 ms 1.83 ms 1.59 ms 1.61 ms
Baseline 0.75 ms 7.87 ms 229.39 ms 352.29 ms 54.17 ms 101.05 ms
Input Meas. 0.68 ms 2.38 ms 62.98 ms 96.70 ms 5.95 ms 11.97 ms
Input Size (MB) 0.05 1.20 36.95 58.96 3.69 7.60
Output Meas. 0.02 ms 0.40 ms 0.50 ms 0.56 ms 0.46 ms 0.60 ms
Attestation 28.97 ms 26.26 ms 27.38 ms 29.26 ms 23.13 ms 31.78 ms
Overhead w/ att 29.67 ms 29.04 ms 90.86 ms 126.52 ms 29.54 ms 44.35 ms
Overhead w/ att (%) 3955.35 369.18 39.61 55.15 54.52 81.86
Signing in s/w 0.18 ms 3.79 ms 4.40 ms 4.84 ms 4.64 ms 4.56 ms
Overhead w/ sgn 0.88 ms 6.56 ms 67.87 ms 102.10 ms 11.05 ms 17.13 ms
Overhead w/ sgn (%) 1.17 0.83 0.30 0.29 0.20 0.17
We describe how to reduce the per-inference overhead. One approach is to batch multiple inputs from a single client and attest them together, spreading the cost of a single attestation or signature across multiple tasks, resulting in overhead similar to accuracy attestation in Table 5. Another approach to reduce cost is to generate a signing keypair during TA initialization, perform a single attestation of the public key, and then sign each inference result for indirect, low-cost attestation. We compare the time taken for signing with a keypair versus invoking hardware attestation. We find that the resulting overhead computed using signing (“Overhead w/ sgn” in Table 5) is drastically less than using attestation (“Overhead w/ att”).
Summary. These findings highlight Laminator’s efficiency, with minimal overhead in comparison with the baseline cost of the computation being attested. The only exception to this is for inference attestation, where the time spent on inference is more comparable to the time spent on attestation; we suggest a mitigation for this by using signing key-pair in software instead of hardware-based attestation. Verification costs are low, involving only certificate checking, and are independent of model size. Hence, Laminator satisfies R1.
6.2. R2: Scalability
Mainstream TEEs provide attestations based on digital signatures that can be repeatedly verified by multiple verifiers. Additionally, an assertion bundle can include attestations from various provers, which are standardized into a common format. This allows Laminator to scale to large numbers of independent provers, without relying on a single entity or TEE type to generate all proofs needed for a verifiable property card. Hence, Laminator satisfies R2.
6.3. R3: Versatility
Laminator’s property definitions are software-based, with hardware only used for isolation and binary attestation, making it easy to extend to new properties. $\mathcal{P}$s can write Python scripts to measure new properties, get them certified, and $\mathcal{V}$s can validate the claims. The attestation and certification formats map directly to YAML property card metadata, so $\mathcal{V}$ requires no modification to support new metrics or ML property cards. Thus, Laminator satisfies R3.
7. Related Work